OpenStudy (anonymous):

A nurse measures the height of a patient as 53 in. Convert this measurement to feet and inches. ? ft ? in

OpenStudy (anonymous):

Can you tell me how many inches are in a foot?

OpenStudy (anonymous):

36

OpenStudy (anonymous):

no sorry 12

OpenStudy (anonymous):

Yep, 12 is correct. So how many times does 12 go into 53?

OpenStudy (anonymous):

636

OpenStudy (anonymous):

You think 12*636 is less than 53?

OpenStudy (anonymous):

4.416

OpenStudy (anonymous):

Yeah, so there are 4 feet and the remainder is how many inches there are. Can you work out the answer now?

OpenStudy (anonymous):

4 feet 42 inches

OpenStudy (anonymous):

The remainder will be between 1 and 11. We know your answer is wrong because 4*12+42=90 inches, not 53.

OpenStudy (mayankdevnani):

@harz360 understood how to solve it ??

OpenStudy (mayankdevnani):

if any questions,please reply !

OpenStudy (anonymous):

no i dont understand it

OpenStudy (mayankdevnani):

don't worry !!

OpenStudy (mayankdevnani):

first tell me how many inches are there in a foot ?

OpenStudy (anonymous):

12

OpenStudy (mayankdevnani):

good !

OpenStudy (mayankdevnani):

we have to measure 53 inches in a foot and we know that 1foot=12 inch

OpenStudy (mayankdevnani):

\[\large \bf 1 ~foot=12~inches\] \[\large \bf 1~inch=\frac{1}{12}~foot\] so, \[\large \bf 53~inches=\frac{1}{12} \times 53=\frac{53}{12}\]

OpenStudy (mayankdevnani):

when we divide it,we get